On a train, 45% of the passengers are women, 30% are men and 25% are children.
If there are 120 more women than men on the train, how many children are on the train?

Answer by jim_thompson5910(35256)

m = # of men
w = # of women
c = # of children
t = total number of passengers on the train

45% of the passengers are women, so
w = 0.45t

30% of the passengers are men, so
m = 0.30t

-------------------------------------------

"there are 120 more women than men on the train" means
w = m+120

Now plug in m = 0.30t and w = 0.45t

w = m+120
w = 0.30t+120
0.45t = 0.30t+120

and isolate t

0.45t = 0.30t+120
0.45t - 0.30t = 120
0.15t = 120
t = 120/0.15
t = 800

There are 800 people on the train

"25% are children", so

c = 25% of 800 = 0.25*800 = 200

There are 200 children on the train